OpenStudy (vespersaviour):

A flag pole is 12 feet tall. Its shadow is 10 feet long. What is the approximate distance from the top of the flag pole to the tip of the pole's shadow? A. 22 feet B. 16 feet C. 11 feet D. 7 feet

OpenStudy (vespersaviour):

Im thinking its either B or D.

OpenStudy (vespersaviour):

12^2 + 10^2 = x 12^2 = 144 10^2 = 100 144 + 100 = 244 \[\sqrt{244} = 15.62049935181331\] which estimated would be B. ^^^thats the way someone taught me to do it.

OpenStudy (xxanntdmxx):

yep your correct, its B.

OpenStudy (vespersaviour):

a^2 + 10^2 = 12^2 a^2 + 100 = 144 a^2 = 144 - 100 a^2 = 44 a = √44 \[\sqrt{44}= 6.6332495807108\] Which is D